OpenStudy (christos):

Can I find the distance between the points P1 = (3, -1) and P2 = (-1, -4) without using a graph?

hartnn (hartnn):

yes, Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (christos):

Yea I will try it right now :D

hartnn (hartnn):

okay :) and tell me what u get ...to verify.

OpenStudy (christos):

I got square root of 21

hartnn (hartnn):

x1- x2 = 3 - (-1) =... ?

OpenStudy (christos):

25 sorry

hartnn (hartnn):

then its correct :)

OpenStudy (christos):

= 5

hartnn (hartnn):

\(\huge \color{red}{\checkmark}\)

OpenStudy (christos):

;p